OpenStudy (anonymous):

sum of arithmetic sequence 6+10+14...

OpenStudy (aravindg):

what is a=? d=? n=?

OpenStudy (anonymous):

I dont really know.All I have is 6+10+14...+86

OpenStudy (anonymous):

see a is the first term so here a=6 d is the difference between each consecutive terms here d=10-6 =4 ( it is either 10-6=4,14-10=4 ... it is constant for arithmetic sequence) n is the number of term her 86 is the nth term so we need to find n first using nthterm= firstterm +(n-1)*(common difference) i.e 86=6+(n-1)*4 80=(n-1)*4 hence n-1=80/4 n-1=20 to find 6+10+14...+86 we have a=6 d=4 n=21 and the last term=86 thus reqd Sum= (n/2)*(firstterm+lastterm)=21/2 *(6+86)=21/2*92=21*46=966 n=20+1 n=21

OpenStudy (aravindg):

@matricked please donot provide direct worked out solutions like that its against openstudy COC .Instead help the user get to the answer .

OpenStudy (anonymous):

@deeva123 If you're asking a question that involves finding the sum of an arithmetic sequence, you must be learning the topic Series and Sequences. If you're unsure of a question, you must reflect on your notes given or provided by your teacher or notes from your maths textbook. Whatever gives you information about sum of Arithmetic Sequences, will help you.